OpenStudy (anonymous):

3x-4y=7 2x-y=3 using the addition method

OpenStudy (anonymous):

3x-4y=7 8x-4y=28 11x=35 x=3.454545...

OpenStudy (anonymous):

how do you get y

OpenStudy (anonymous):

time (-4) 2x-y=3 you get -8x+4y=-12 - 8x+4y=-12 3x-4y=7 ADD--------- -5x=-5 x =1 substitution x =1 find y 3x-4y=7 3(1)-4y=7 3-4y=7 -4y=7-3 y = 4/-4 y =-1

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

ywc,if u like my answer click good answer

OpenStudy (anonymous):

ty